OpenStudy (anonymous):

@Dumbcow Complete the equation that is related to the given equation. m – 18.2 = –58 m = –58 +

OpenStudy (anonymous):

@sunnyshores @phi

OpenStudy (phi):

m – 18.2 = –58 add +18.2 to both sides

OpenStudy (anonymous):

Firstly you need to start by simplifying M + -18.2 = -58 Then of course record the terms: -18.2 + M = -58 Solving -18.2 + M = -58 Solving for variable 'M'. Move all terms containing M to the left, all other terms to the right. Add '18.2' to each side of the equation. -18.2 + 18.2 + M = -58 + 18.2 Next combine like terms: -18.2 + 18.2 = 0.0 0.0 + M = -58 + 18.2 M = -58 + 18.2 Combine like terms: -58 + 18.2 = -39.8 M = -39.8 Lastly you need to simplify again. M = -39.8